OpenStudy (anonymous):

Identify the rule of the function table. Input x 2,4,10 output y 14,28,70? Choices A. Y= 7x B. Y=x+12 C. y= 2x +10. D. Y= 12x.

OpenStudy (here_to_help15):

This ^

OpenStudy (anonymous):

?

OpenStudy (anonymous):

please help

jimthompson5910 (jim_thompson5910):

One input/output pair is x = 2 and y = 14 Plug x = 2 into each equation. Then evaluate and simplify. Which equation produces y = 14?

jimthompson5910 (jim_thompson5910):

Do the same for x = 4, y = 28 and x = 10, y = 70

OpenStudy (anonymous):

y=7x by trial and error.

OpenStudy (anonymous):

1:7, 70/10 = 7 28/4 = 7 14/2 = 7 there are seven of the value of Y for every X so your answer would be A

OpenStudy (anonymous):

thanks for explaining